1.ph meter gives an electrical response proportional to the concentration of hydrogen ion. to convert this








response into ph, glass electrode must be calibrated against standard solution of known h+








concentration.








which of the following standard solution(s) is/are likely to be used?








a) ph 4.0 buffer b) ph 7.0 buffer c) ph 14.0 buffer and d) all of the above

1
Expert's answer
2021-11-26T05:57:01-0500

pH of solution needed = 7.00

pKa of H2PO4- = 7.20

we have,

(H2PO4-) + (HPO4^2-) = 0.100 M x 1 L = 0.100 mol

Using Hendersen-Hasselbalck equation,

pH = pKa + log(base/acid)

7.00 = 7.20 + log(HPO4^2-/H2PO4-)

(HPO4^2-) = 0.631(H2PO4-)

feeding in the above equation,

(H2PO4-) + 0.631(H2PO4-) = 0.100 mol

(H2PO4-) mol = 0.100/1.631 = 0.061 mol

mass NaH2PO4.H2O needed = 0.061 mol x 138.01 g/mol = 8.5 g

(HPO4^2-) mol = 0.100 - 0.061 = 0.039 mol

mass Na2HPO4 needed = 0.039 mol x 141.98 g/mol = 5.5